### v1

# Enhancing Time Series Forecasting via Multi-Level Text Alignment with LLMs

Abstract
The adaptation of large language models (LLMs) to time series forecasting poses unique challenges, as time series data is continuous in nature, while LLMs operate on discrete tokens. Despite the success of LLMs in natural language processing (NLP) and other structured domains, aligning time series data with language-based representations while maintaining both predictive accuracy and interpretability remains a significant hurdle. Existing methods have attempted to reprogram time series data into text-based forms, but these often fall short in delivering meaningful, interpretable results. In this paper, we propose a multi-level text alignment framework for time series forecasting using LLMs that not only improves prediction accuracy but also enhances the interpretability of time series representations. Our method decomposes time series into trend, seasonal, and residual components, which are then reprogrammed into component-specific text representations. We introduce a multi-level alignment mechanism, where component-specific embeddings are aligned with pre-trained word tokens, enabling more interpretable forecasts. Experiments on multiple datasets demonstrate that our method outperforms state-of-the-art models in accuracy while providing good interpretability.
keywords:
Time Series alignment LLMs.
1 Introduction
Time series forecasting, which involves predicting future values based on historical data, has numerous practical applications, such as demand planning, inventory optimization, energy load forecasting, and climate modeling [7, 13, 14, 6]. Traditionally, these tasks demand substantial domain expertise and carefully designed models tailored to specific datasets. However, recent advancements in pre-trained large language models (LLMs), such as GPT-4 [1] and LLaMA [23], have achieved remarkable success in natural language processing (NLP) and demonstrated potential in handling complex, structured domains. This raises a compelling question: how can these powerful pre-trained LLMs be effectively adapted for time series forecasting?
LLMs, trained on vast and diverse text corpora, provide a powerful foundation for various downstream tasks, requiring only minimal task-specific prompt engineering or fine-tuning. This flexibility has sparked a growing interest in leveraging LLMs for time series analysis. For example, methods like Promptcast [28] and LLMTime [8] reformulate numerical inputs and outputs into prompts, treating time series forecasting as a sentence-to-sentence task, which enables the direct application of LLMs. Meanwhile, approaches like TEMPO [2] and GPT4TS [33] take a different route by fine-tuning pre-trained LLMs, modifying components such as the Add&Norm layers and positional embeddings, further demonstrating LLMs’ adaptability for time series forecasting.
[…]
Figure 1: Cross-modality time series embeddings of $(a)$ contrast learning of text-prototype-aligned time series embeddings, $(b)$ text prototypes reprogramming, $(c)$semantic informed prompt and and $(d)$ proposed anchors alignment of our multi-level alignment.
In this paper, we address the challenge of interpretability in LLM-based time series forecasting by developing an interpretable multi-level text alignment framework while preserving the backbone model. Our approach consists of two key principles for effective time series representation learning: (a) modeling specific time series components such as trend, seasonality, and residuals, and (b) deriving interpretable explanations from the inherent properties of time series data through multi-level text alignment. Specifically, we decompose the time series input into three additive components—trend, seasonality, and residuals—using locally weighted scatterplot smoothing (LOESS) [5]. These components are then reprogrammed into component-specific anchors that better align with the language capabilities of LLMs in Figure 1(d). Additionally, we employ component-specific prompts to guide the generation of learnable continuous vector representations that encode temporal knowledge of each component.
In summary, the main contributions of this paper are as follows: (1) We propose an interpretable multi-level text alignment framework for time series forecasting using LLMs, while keeping the backbone model unchanged. (2) Our method leverages this multi-level alignment to map decomposed time series components—trend, seasonality, and residuals—into distinctive, informative joint representations. The aligned trend-specific anchors enhance the interpretability of LLMs, while the aligned seasonality and residual prototypes improve the overall representation of the input time series. (3) Experimental results on multiple datasets validate the superiority of our model over state-of-the-art approaches, highlighting the effectiveness of interpretable multi-level text alignment.
2 Related work
2.1 Pre-trained Large Language Models for Time series.
The recent advancements in Large Language Models (LLMs) have opened up new opportunities for time series modeling. LLMs like T5 [21], GPT-2 [20], GPT-4 [1], and LLaMA [23] have demonstrated impressive capabilities in understanding complex dependencies in heterogeneous textual data and generating meaningful outputs. Recently, there has been growing interest in exploring how to transfer the knowledge embedded in these pre-trained LLMs to the time series domain [11, 9]. For instance, [28] converts time series data into text sequences, achieving promising results. Other works, such as [33, 8], tokenize time series data into overlapping patches and strategically fine-tune LLMs for time series forecasting tasks. Similarly, recent works such as [2, 19] decompose time series data and use retrieval-based prompts to enhance fine-tuning of pre-trained LLMs. However, these approaches often fall short of delivering interpretable results and tend to treat time series as mere sequences of tokens, overlooking their inherent temporal structures. Converting numerical data to text without sufficient alignment to temporal dynamics can lead to inaccurate predictions and a lack of transparency in the model’s decision-making process, especially for multivariate time series. Our work introduces an interpretable multi-level text alignment framework to align time series components with anchors while keeping pre-trained LLMs intact.
2.2 Time Series Aligned Embeddings
[…]
Figure 2: The architecture of the proposed multi-level aligned embeddings begins with the decomposition of the input time series into three components: trend, seasonal, and residual. These components, tokenized and embedded, are reprogrammed using anchors and condensed text prototypes to align time series data with word tokens. Component-specific prefixed prompts are added to guide the transformation of input patches. Outputs from the LLM are projected, de-normalized, and summed to generate the final prediction.
3 Methodology
Our approach focuses on enhancing the interpretability of large language models (LLMs) for time series data through multi-level aligned embeddings. As illustrated in Figure 2, the proposed framework consists of four core modules: (1) time series input decomposition, (2) multi-level text alignment, (3) component-specific prompts, and (4) output projection. The process begins by partitioning a multivariate time series into $N$ univariate time series, each processed independently. The $i$-th series, denoted as $\mathbf{X}^{(i)}\in\mathbb{R}^{1\times L}$, undergoes a series of steps including decomposition, normalization, patching, and embedding before being aligned with anchor points and text prototypes. To enhance the LLM’s reasoning capability on time series data, we introduce component-specific prompts along with the aligned embeddings, enabling the model to generate meaningful output representations. These representations are then projected through an output linear layer to produce the final forecasts, $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$. With the primary objective of improving interpretability, we utilize GPT-2 [20], employing its first six layers as the backbone model for time series forecasting without fine-tuning the foundational model.
3.1 Problem statement
Given the observed values over the previous $L$ timestamps, the task of multivariate time series forecasting is to predict the values for the next $H$ timestamps. Formally, this can be represented as:
$\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}=F\left(\mathbf{x}_{t-L}^{(i)},\dots,\mathbf{x}_{t-1}^{(i)};\mathbf{V}^{(i)}\right),$ (1)
where $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$ is the vector of $H$-step prediction from timestamp $t$ of channel $i$ corresponding to the $i$-th feature. Given the historical values $\mathbf{x}_{t-L}^{(i)},\dots,\\ \mathbf{x}_{t-1}^{(i)}$, a large language model $F$ uses prompt $\mathbf{V}^{(i)}$ to make these predictions. Leveraging the strong reasoning capabilities of pre-trained large language models, we aim to align time series data with text to enable LLMs to interpret the input series and accurately forecast the $H$ future steps, with the overall objective of minimizing the mean square errors between the ground truths and predictions, expressed as:
$\frac{1}{H}\sum_{h=1}^{H}\|\hat{\mathbf{x}}_{t-1+h}^{(i)}-\mathbf{x}_{t-1+h}^{(i)}\|^{2}.$ (2)
3.2 Time series input decomposition
For time series data, decomposing complex inputs into meaningful components such as trend, seasonal, and residual elements can help optimally extract valuable information. In this paper, given the input $X\in\mathbb{R}^{N\times L}$, where $N$ is the feature size and $L$ is the length of the time series, the additive decomposition can be represented as:
$X^{(i)}=X^{(i)}_{T}+X^{(i)}_{S}+X^{(i)}_{R},$ (3)
where $i$ refers to the feature index for multivariate time series input. The trend component $X_{T}\in\mathbb{R}^{N\times L}$, capturing the underlying long-term patterns in the data, is expressed as $X_{T}=\frac{1}{m}\sum_{j=-k}^{k}X_{t+j}$, where $m=2k+1$ and $k$ is the averaging step size. The seasonal component $X_{S}\in\mathbb{R}^{N\times L}$ reflects the repeating short-term cycles and can be estimated after removing the trend. The residual component $X_{R}\in\mathbb{R}^{N\times L}$ represents the remainder of the data once the trend and seasonal elements have been extracted. There are multiple methods available for performing additive seasonal-trend decomposition. One common approach is the classical additive seasonal-trend decomposition, which first extracts the long-term trend using moving averages. The seasonal component is then estimated by averaging the detrended time series, and the residual is obtained by subtracting the estimated trend and seasonal components. Another widely used method is the Seasonal-Trend decomposition using Loess (STL) [5]. The choice of decomposition method in this paper is determined based on validation results.
Following the approach outlined in [17], we patch the decomposed components of the time series. Specifically, for the $i$-th normalized trend component, we obtain the patched token $\textbf{P}^{(i)}_{T}\in\mathbb{R}^{L_{P}\times K}$, where $L_{P}$ represents the patch length and $K=\lfloor\frac{(L-L_{P})}{s}\rfloor+2$ denotes the number of patches, with $s$ is the stride. Similarly, we apply this patching process to the seasonal and residual components, obtaining patched tokens $\textbf{P}^{(i)}_{S}$ and $\textbf{P}^{(i)}_{R}$, respectively. These patched tokens are then fed into the multi-level text alignment module to produce aligned time series embeddings.
3.3 Multi-level text alignment
Here we reprogram patch embeddings into the LLMs’ pre-training data representation space to align the modalities of time series and natural language to activate the backbone’s time series understanding and reasoning capabilities. Naively, we can align the token embedding of time series and text using similarity estimation. Although time series tokens lack text annotation, we can place their embedding near typical text descriptions of time series. Thus, it is intuitively expected that various time series tokens can represent various descriptive words such as up, down, stable, and so on. However, the pre-trained word token embedding space is vast and dense, and the selection of text prototypes (patterns) is often highly relaxed, sometimes even involving random words unrelated to time series or clusters of pre-trained word tokens [22, 19, 10], which leads to poor interpretability.
In this work, we propose multi-level text alignment to enhance the interpretability of LLMs on time series forecasting. We first decompose the time series into trend, seasonal, and residual and align the trend $X_{T}$ with selected trend-specific anchors $\mathbb{W}_{trend}$. However, accurately defining anchors for the seasonal and residual components is challenging. To address this, we reprogram seasonal $X_{S}$ and residual $X_{R}$ using pre-trained word embeddings $\textbf{E}\in\mathbb{R}^{V\times D}$ in the backbone, where $V$ is the vocabulary size, $D$ is the hidden dimension of the pre-trained LLM. Directly leveraging E will result in large and potentially dense reprogramming space. We adapt linearly probing E. The text prototypes of seasonal and residual denoted as $\textbf{E}^{{}^{\prime}}_{seasonal}\in\mathbb{R}^{V^{{}^{\prime}}_{seasonal}\times D}$ and $\textbf{E}^{{}^{\prime}}_{residual}\in\mathbb{R}^{V^{{}^{\prime}}_{residual}\times D}$, where $V^{{}^{\prime}}_{seasonal}<V^{{}^{\prime}}_{residual}\ll V$ because the residual is more inconsistent and variable compared to the seasonal.
As illustrated in the top-right of Figure 2, our multi-level text alignment aims to give a connection between anchors and trend patches. The selected anchors are sparse. We reprogram seasonal and residual with text prototypes connecting time series patches with a more dense reprogramming space. To realize this, we employ a multi-head cross-attention layer for each component. Specifically, for $i$-th input feature, we define query matrices $\mathbf{Q}_{T}^{(i)}=\textbf{P}_{T}^{(i)}\textbf{W}^{Q(i)}_{T}$, key matrices $\mathbf{K}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{K(i)}_{T}$, value matrices $\mathbf{V}_{T}^{(i)}=\mathbb{W}_{trend}\textbf{W}^{V(i)}_{T}$ for trend; query matrices $\mathbf{Q}_{S}^{(i)}=\textbf{P}_{S}\textbf{W}^{Q(i)}_{S}$, key matrices $\mathbf{K}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{K(i)}_{S}$, value matrices $\mathbf{V}_{S}^{(i)}=\textbf{E}^{{}^{\prime}}_{seasonal}\textbf{W}^{V(i)}_{S}$ for seasonal; query matrices $\mathbf{Q}_{R}^{(i)}=\textbf{P}_{R}\textbf{W}^{Q(i)}_{R}$, key matrices $\mathbf{K}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{K(i)}_{R}$, value matrices $\mathbf{V}_{R}^{(i)}=\textbf{E}^{{}^{\prime}}_{residual}\textbf{W}^{V(i)}_{R}$ for residual. Through multi-head attention, we reprogram each time series component. For example, the trend after multi-head attention is defined as:
$\mathbf{Z}_{T}^{(i)}=\text{ATTENTION}(\mathbf{Q}_{T}^{(i)},\mathbf{K}_{T}^{(i)},\mathbf{V}_{T}^{(i)})=\text{SOFTMAX}\left(\frac{\mathbf{Q}_{T}^{(i)}\mathbf{K}_{T}^{(i)\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{T}^{(i)},$ (4)
where $d_{k}$ is the dimension of each head in the multi-head attention module. After the multi-head attention step, each component is linearly projected to align the hidden dimensions with the backbone model.
3.4 Component-specific prompts
Prompting techniques have proven highly effective across various tasks by leveraging task-specific knowledge encoded in prompts. This success stems from prompts providing a structured framework that aligns the model’s output with desired objectives, improving accuracy and coherence. However, directly translating time series into natural language poses challenges, complicating the creation of instruction-following datasets and effective on-the-fly prompting [27]. Recent advances show that prompts can enrich input context and guide the transformation of reprogrammed time series patches [10]. To leverage the semantic information of time series components, we propose a component-specific prefix prompting strategy. This includes three elements: dataset context, input statistical features, and component-specific task instructions for trend, seasonal, and residual components. For instance, the task description ’forecast the next 96 steps given the previous 512 steps [trend, seasonal, residual]’ serves as a template for our task instructions, which are then concatenated with the corresponding component data.
3.5 Output projection
After packing and forwarding the component-specific prompts and embeddings through the frozen backbone LLM, we retain the embedding for each component and apply a linear projection to the output representation. By denormalizing and summing these representations, we derive the final forecasts $\hat{\mathbf{x}}_{t}^{(i)},\dots,\hat{\mathbf{x}}_{t+H-1}^{(i)}$.
4 Experiments
In our experiments, the proposed method outperforms state-of-the-art forecasting approaches across various benchmarks, including long-term, short-term, and few-shot forecasting. For a fair comparison, we follow the configurations outlined in [25] across all baselines, utilizing a unified evaluation pipeline [^1] https://github.com/thuml/Time-Series-Library. Our code will be made available on GitHub upon the acceptance of the paper.
Baselines. We compare with the SOTA time series models and cite their performance from [33, 4] if applicable. The SOTA includes a set of Transformer-based methods, i.e., PatchTST [17], ETSformer [24], Non-Stationary Transformer [16], FEDformer [32], Autoformer [26], Informer [31], and Reformer [12]. We also select a set of non-transformer-based techniques, i.e., DLinear [29], TimesNet [25], N-BEATS [18], and LightTS [30]. Finally, four methods are based on LLMs, i.e., TimeLLM [10], LLM4TS [4], GPT4TS [33], and LLMTime [8]. Aligned with the GPT4TS configuration [33], we utilize only the first 6 layers of the 12-layer GPT-2 base [20] as the backbone model of ours and TimeLLM.
4.1 Long-term Forecasting
Setup. For long-term forecasting, we evaluate on ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity(ECL), and Traffic, which have been widely adopted as benchmarking datasets for long-term forecasting works [25]. The input time series length $L$ is set as 512, and we evaluate across four prediction horizons: $H\in\{96,192,336,720\}$. The evaluation metrics include mean square error (MSE) and mean absolute error (MAE).
Results. Table 1 presents the performance of various time series forecasting models on MSE and MAE metrics across different prediction horizons on multiple benchmarks. Our proposed model consistently outperforms existing baselines, demonstrating superior performance on average across most datasets and prediction lengths. This highlights the applicability of multi-level text alignment. Notably, our comparison with TimeLLM—a recent work leveraging text prototype reprogramming to align time series with text tokens—is significant. Specifically, our model achieves substantial improvements on the Weather and ETTm1 datasets, exceeding the best-performing LLM-based model, LLM4TS, by 23.3% and 26.8%, respectively, in terms of MSE. Additionally, it records the lowest error rates across numerous dataset-prediction length configurations. These results suggest that integrating LLMs with multi-level text alignment can significantly enhance the accuracy of long-term time series forecasting.
Table 1: Long-term forecasting results for {96, 192, 336, 720} horizons. Lower values indicate better performance. bold: the best, underline: second best.
Methods Ours Time-LLM LLM4TS GPT4TS DLinear PatchTST
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.355 0.404 0.384 0.407 0.371 0.394 0.376 0.397 0.375 0.399 0.370 0.399
192 0.426 0.445 0.423 0.434 0.403 0.412 0.416 0.418 0.405 0.416 0.413 0.421
336 0.434 0.449 0.435 0.447 0.420 0.422 0.442 0.433 0.439 0.443 0.422 0.436
720 0.480 0.493 0.439 0.463 0.422 0.444 0.477 0.456 0.472 0.490 0.447 0.466
Avg. 0.424 0.448 0.420 0.438 0.404 0.418 0.428 0.426 0.422 0.437 0.413 0.435
ETTh2 96 0.260 0.336 0.295 0.355 0.269 0.332 0.285 0.342 0.289 0.353 0.274 0.336
192 0.333 0.375 0.376 0.410 0.328 0.377 0.354 0.389 0.383 0.418 0.339 0.379
336 0.369 0.408 0.376 0.412 0.353 0.396 0.373 0.407 0.448 0.465 0.329 0.380
720 0.444 0.455 0.410 0.442 0.383 0.425 0.406 0.441 0.605 0.551 0.379 0.422
Avg. 0.378 0.408 0.364 0.403 0.333 0.383 0.355 0.394 0.431 0.446 0.330 0.379
ETTm1 96 0.117 0.232 0.297 0.349 0.285 0.343 0.292 0.346 0.299 0.343 0.290 0.342
192 0.198 0.298 0.336 0.373 0.324 0.366 0.332 0.372 0.335 0.365 0.332 0.369
336 0.301 0.360 0.362 0.390 0.353 0.385 0.366 0.394 0.369 0.386 0.366 0.392
720 0.389 0.411 0.410 0.421 0.408 0.419 0.417 0.421 0.425 0.421 0.416 0.425
Avg. 0.251 0.325 0.351 0.383 0.343 0.378 0.352 0.383 0.357 0.378 0.351 0.380
ETTm2 96 0.095 0.200 0.177 0.264 0.165 0.254 0.173 0.262 0.167 0.269 0.165 0.255
192 0.174 0.263 0.253 0.312 0.220 0.292 0.229 0.301 0.224 0.303 0.220 0.292
336 0.243 0.313 0.285 0.345 0.268 0.326 0.286 0.341 0.281 0.342 0.274 0.329
720 0.343 0.380 0.366 0.390 0.350 0.380 0.378 0.401 0.297 0.421 0.362 0.385
Avg. 0.214 0.289 0.270 0.328 0.251 0.313 0.267 0.326 0.267 0.333 0.255 0.315
Weather 96 0.059 0.125 0.158 0.210 0.147 0.196 0.162 0.212 0.176 0.237 0.149 0.198
192 0.115 0.188 0.191 0.240 0.191 0.238 0.204 0.248 0.220 0.282 0.194 0.241
336 0.211 0.263 0.247 0.284 0.241 0.277 0.254 0.286 0.265 0.319 0.245 0.282
720 0.299 0.327 0.319 0.334 0.313 0.329 0.326 0.337 0.333 0.362 0.314 0.334
Avg. 0.171 0.226 0.229 0.267 0.223 0.260 0.237 0.271 0.248 0.300 0.225 0.264
ECL 96 0.116 0.221 0.137 0.237 0.128 0.223 0.139 0.238 0.140 0.237 0.129 0.222
192 0.145 0.250 0.150 0.249 0.146 0.240 0.153 0.251 0.153 0.249 0.150 0.240
336 0.167 0.271 0.168 0.266 0.163 0.258 0.169 0.266 0.169 0.267 0.163 0.259
720 0.209 0.307 0.203 0.293 0.200 0.292 0.206 0.297 0.203 0.301 0.197 0.290
Avg. 0.159 0.262 0.164 0.261 0.159 0.253 0.167 0.263 0.166 0.263 0.161 0.252
Traffic 96 0.255 0.229 0.380 0.277 0.372 0.259 0.388 0.282 0.410 0.282 0.360 0.249
192 0.332 0.258 0.399 0.288 0.391 0.265 0.407 0.290 0.423 0.287 0.379 0.256
336 0.370 0.273 0.408 0.290 0.405 0.275 0.412 0.294 0.436 0.296 0.392 0.264
720 0.428 0.301 0.445 0.308 0.437 0.292 0.450 0.312 0.466 0.315 0.432 0.286
Avg. 0.346 0.265 0.408 0.290 0.401 0.273 0.414 0.294 0.433 0.295 0.390 0.263
4.2 few-shot forecasting
Setups. LLMs have recently shown impressive few-shot learning capabilities [15]. To evaluate performance in the few-shot forecasting setting, we follow the experimental setup outlined in [33], allowing us to assess whether the model can generate accurate forecasts with limited training data. In these experiments, we use only the first 10% of the training data.
Results. The brief 10% few-shot learning results in Table 2 and full results in Table 5 demonstrate that our model significantly outperforms all baseline methods across most datasets. We attribute this success to the effective knowledge activation achieved through multi-level text alignment. Specifically, our model improves few-shot learning performance on the Weather and Traffic datasets by 11.9% and 5%. When trained with only 10% of the data, LLM-based methods substantially outperform other baselines, which are trained from scratch and thus limited by the smaller training set. In contrast, LLM-based models can leverage pre-trained knowledge and align it with time series embeddings to enhance representation.
Table 2: Few-shot learning on 10% training data. All results are averaged from four different forecasting horizons: $H\in\{96,192,336,720\}$. Lower values indicate better performance.
Methods Ours TimeLLM LLM4TS GPT4TS DLinear PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 0.397 0.431 0.446 0.464 0.366 0.407 0.397 0.421 0.605 0.538 0.415 0.431
ETTm2 0.262 0.324 0.292 0.343 0.276 0.324 0.293 0.335 0.316 0.368 0.296 0.343
Weather 0.207 0.263 0.359 0.275 0.235 0.270 0.238 0.275 0.241 0.283 0.242 0.279
ECL 0.190 0.288 0.182 0.277 0.172 0.264 0.176 0.269 0.180 0.280 0.180 0.273
Traffic 0.409 0.310 0.438 0.312 0.432 0.303 0.440 0.310 0.447 0.313 0.430 0.305
Table 3: Zero-shot learning results on ETT datasets. Lower values indicate better performance. bold: the best, underline: second best.
Methods Ours Time-LLM GPT4TS LLMTime PatchTST DLinear
Datasets MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 $\rightarrow$ ETTh2 0.346 0.396 0.354 0.400 0.406 0.422 0.992 0.708 0.380 0.405 0.493 0.488
ETTh1 $\rightarrow$ ETTm2 0.294 0.357 0.310 0.363 0.325 0.363 1.867 0.869 0.314 0.360 0.415 0.452
ETTh2 $\rightarrow$ ETTm2 0.276 0.345 0.303 0.356 0.335 0.370 1.867 0.869 0.325 0.365 0.328 0.386
ETTm1 $\rightarrow$ ETTm2 0.217 0.284 0.275 0.325 0.313 0.348 1.867 0.869 0.296 0.334 0.335 0.389
ETTm2 $\rightarrow$ ETTm1 0.562 0.478 0.501 0.453 0.769 0.567 1.933 0.984 0.568 0.492 0.649 0.537
4.3 Zero-shot forecasting
Setups. Beyond few-shot learning, LLMs also show promise as effective zero-shot learners. In this section, we evaluate the zero-shot learning capabilities of the multi-level text-aligned LLM. Specifically, we assess how well the model performs on one dataset after being optimized on another. Similar to the few-shot learning setup, we use the long-term forecasting protocol and evaluate various cross-domain scenarios utilizing the ETT datasets.
Results. The brief results are presented in Table 3, with full results in Table 6. Our model demonstrates performance that is comparable to or surpasses other baselines. In data-scarce scenarios, our model significantly outperforms other LLM-based models, consistently providing better forecasts. Both our model and TimeLLM [10] outperform traditional baselines, likely due to cross-modality alignment, which more effectively activates LLMs’ knowledge transfer and reasoning capabilities for time series tasks. Additionally, our multi-level aligned embeddings better align language cues with temporal components of time series, enabling superior zero-shot forecasting performance compared to TimeLLM.
4.4 Model analysis
Table 4: Comparison of different variants for long-term and few-shot forecasting.
Variant Long-term Forecasting Few-shot Forecasting
ETTm1-96 ETThm1-192 ETTm1-96 ETTm1-192
Default GPT-2 (6) 0.117 0.198 0.360 0.429
A.1 w/o alignment 0.262 0.347 0.571 0.583
B.1 only trend alignment 0.184 0.283 0.476 0.578
B.2 only seasonal alignment 0.127 0.212 0.367 0.432
B.3 only residual alignment 0.171 0.229 0.433 0.506
C.1 noise anchors 0.134 0.214 0.424 0.464
C.2 synonymous anchors 0.119 0.202 0.366 0.434
D.1 w/o component-specific instruction 0.125 0.205 0.408 0.461
D.2 w/o domain features 0.118 0.199 0.371 0.440
Multi-level text alignment variants. Our results in Table 4 show that removing component alignment or prefixed prompts negatively impacts knowledge transfer during LLM reprogramming for effective time series forecasting. Specifically, without alignment (A.1), we observe a significant average performance drop of 75.4% across standard and few-shot forecasting tasks. We also examine the effect of aligning only two components to assess whether aligning just one component is sufficient in our multi-level alignment strategy. Retaining only seasonal alignment (B.2) achieves the best performance, though it still results in an average MSE increase of 4.5% across all scenarios. In contrast, keeping only trend alignment significantly degrades performance, with over 32.2% performance loss in both standard and few-shot tasks. Furthermore, altering the selection of anchors for trend alignment (C.1, C.2) increases MSE by over 14.5% when using noise anchors. Expanding the anchor selection with synonymous words produces results comparable to the default, with less than a 2% variation in MSE. Finally, removing component-specific instruction (D.1) and domain features (D.2) results in MSE increases of 7.7% and 1.7%, respectively.
Multi-level text alignment interpretation. We present a case study on ETTm1 using non-overlapping patching, where the patch stride is the same as the patch length, with different selected anchors shown in Figure 3. In our main experiments, the original anchors are “increase decrease upward downward linear exponential drift stable volatile stationary persistent rapid”. The attention map illustrates the optimized attention scores between input trend patches and aligned anchors. These matched anchors serve as textual shapelets for the time series tokens. Specifically, subplot (a) displays the optimized attention scores for synonymous anchors, which are consistent. The highlighted anchors, “rise”, “increase”, “climb”, “grow”, and “expand”, are associated with upward trend patches. In contrast, subplot (b) shows no highlighted anchor when all trend patches are aligned with noise words unrelated to time series trends. This case study demonstrates that aligned anchors effectively summarize the textual shapelets of the input trend patches.
Figure 3: A showcase of visualization of multi-level alignment interpretation.
Although multi-level alignment in both seasonal and residual components can provide visual interpretations, visualizing these alignments is challenging. Since the input patches are aligned with text prototypes learned from a large and dense pre-trained word embedding space, more tools are needed to present a better visualization across two optimized layers. Moreover, the trend component is the most interpretable and semantically clear of the three components, in contrast to the noisy residual and the seasonal component, which lacks textual semantics. Our model efficiencies in terms of parameters, memory, and speed are comparable to TimeLLM [10] with only two additional lightweight aligned embedding layers for integrating trend and seasonal components.
5 Conclusion and future work
We propose a multi-level text alignment framework with pre-trained language models for time series forecasting. Our multi-level aligned embeddings enhance the LLM’s interpretability and forecasting performance by aligning time series components with anchors and text prototypes. Our results demonstrate that time series tokens aligned with anchors provide a clearer and more intuitive interpretation of similar time series trends. Future research should focus on optimizing the alignment module for selected anchors and time series tokens, and work toward developing multimodal models capable of joint reasoning across time series, natural language, and other modalities.
Appendix A Appendix
Our full results in few-shot forecasting tasks are detailed in Table 5. With the scope of 10% few-shot learning, our model’s secure SOTA performance in 13 out of 35 cases, spanning seven different time series benchmarks. Moreover, our model only lose to LLM4TS, which is neither interpretable nor light-weight.
Table 5: Full few-shot results on 10% training data.
Methods Ours Time-LLM LLM4TS GPT4TS DLinear PatchTST
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.629 0.548 0.530 0.492 0.417 0.432 0.458 0.456 0.492 0.495 0.516 0.485
192 0.720 0.534 0.671 0.546 0.469 0.468 0.570 0.516 0.565 0.538 0.598 0.524
336 0.893 0.622 0.907 0.639 0.505 0.499 0.608 0.535 0.721 0.622 0.657 0.550
720 1.210 0.772 0.917 0.647 0.708 0.572 0.725 0.591 0.986 0.743 0.762 0.610
Avg. 0.865 0.619 0.756 0.581 0.525 0.493 0.590 0.525 0.691 0.600 0.633 0.542
ETTh2 96 0.330 0.383 0.349 0.418 0.282 0.351 0.331 0.374 0.357 0.411 0.353 0.389
192 0.357 0.408 0.406 0.428 0.364 0.400 0.402 0.411 0.569 0.519 0.403 0.414
336 0.406 0.439 0.488 0.489 0.374 0.416 0.406 0.433 0.671 0.572 0.426 0.441
720 0.498 0.496 0.540 0.520 0.445 0.461 0.449 0.464 0.824 0.648 0.477 0.480
Avg. 0.397 0.431 0.446 0.464 0.366 0.407 0.397 0.421 0.605 0.538 0.415 0.431
ETTm1 96 0.360 0.389 0.297 0.349 0.360 0.388 0.390 0.404 0.352 0.392 0.410 0.419
192 0.429 0.431 0.336 0.373 0.386 0.401 0.429 0.423 0.382 0.412 0.437 0.434
336 0.446 0.465 0.362 0.390 0.415 0.417 0.469 0.439 0.419 0.434 0.476 0.454
720 0.489 0.495 0.410 0.421 0.470 0.445 0.569 0.498 0.490 0.477 0.681 0.556
Avg. 0.431 0.445 0.351 0.383 0.402 0.457 0.464 0.441 0.411 0.429 0.501 0.466
ETTm2 96 0.126 0.231 0.192 0.276 0.184 0.265 0.188 0.269 0.213 0.303 0.191 0.274
192 0.223 0.300 0.266 0.320 0.240 0.301 0.251 0.309 0.278 0.345 0.252 0.317
336 0.290 0.345 0.317 0.356 0.294 0.337 0.307 0.346 0.338 0.385 0.306 0.353
720 0.412 0.420 0.418 0.420 0.386 0.393 0.426 0.417 0.436 0.440 0.433 0.427
Avg. 0.262 0.324 0.292 0.343 0.276 0.324 0.293 0.335 0.316 0.368 0.296 0.343
Weather 96 0.102 0.177 0.164 0.220 0.158 0.207 0.163 0.215 0.171 0.224 0.165 0.215
192 0.164 0.234 0.215 0.258 0.204 0.249 0.210 0.254 0.215 0.263 0.210 0.257
336 0.230 0.281 0.259 0.294 0.254 0.288 0.256 0.292 0.258 0.299 0.259 0.297
720 0.334 0.363 0.319 0.326 0.322 0.336 0.321 0.339 0.320 0.346 0.332 0.346
Avg. 0.207 0.263 0.359 0.275 0.235 0.270 0.238 0.275 0.241 0.283 0.242 0.279
ECL 96 0.144 0.250 0.145 0.246 0.135 0.231 0.139 0.237 0.150 0.253 0.140 0.238
192 0.167 0.271 0.160 0.259 0.152 0.246 0.156 0.252 0.164 0.264 0.160 0.255
336 0.194 0.294 0.182 0.278 0.173 0.267 0.175 0.270 0.181 0.282 0.180 0.276
720 0.255 0.340 0.239 0.324 0.229 0.312 0.233 0.317 0.223 0.321 0.241 0.323
Avg. 0.190 0.288 0.182 0.277 0.172 0.264 0.176 0.269 0.180 0.280 0.180 0.273
Traffic 96 0.347 0.292 0.416 0.295 0.402 0.288 0.414 0.297 0.419 0.298 0.403 0.289
192 0.398 0.305 0.424 0.306 0.416 0.294 0.426 0.301 0.434 0.305 0.415 0.296
336 0.427 0.313 0.435 0.314 0.429 0.302 0.434 0.303 0.449 0.313 0.426 0.304
720 0.466 0.330 0.476 0.331 0.480 0.326 0.487 0.337 0.484 0.336 0.474 0.331
Avg. 0.409 0.310 0.438 0.312 0.432 0.303 0.440 0.310 0.447 0.313 0.430 0.305
[…]
Table 6: Full zero-shot learning results on ETT datasets.
Methods Ours Time-LLM GPT4TS LLMTime PatchTST DLinear
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 $\rightarrow$ ETTh2 96 0.263 0.337 0.264 0.340 0.335 0.374 0.510 0.576 0.304 0.350 0.347 0.400
192 0.315 0.374 0.332 0.376 0.412 0.417 0.523 0.586 0.386 0.400 0.447 0.460
336 0.371 0.414 0.395 0.424 0.441 0.444 0.640 0.637 0.414 0.428 0.515 0.505
720 0.435 0.462 0.423 0.459 0.438 0.452 2.296 1.034 0.419 0.443 0.665 0.589
Avg. 0.346 0.396 0.354 0.400 0.406 0.422 0.992 0.708 0.380 0.405 0.493 0.488
ETTh1 $\rightarrow$ ETTm2 96 0.191 0.296 0.224 0.311 0.236 0.315 0.646 0.563 0.215 0.304 0.255 0.357
192 0.259 0.338 0.270 0.339 0.287 0.342 0.934 0.654 0.275 0.339 0.338 0.413
336 0.317 0.370 0.336 0.378 0.341 0.374 1.157 0.728 0.334 0.373 0.425 0.465
720 0.409 0.424 0.410 0.422 0.435 0.422 4.730 1.531 0.431 0.424 0.640 0.573
Avg. 0.294 0.357 0.310 0.363 0.325 0.363 1.867 0.869 0.314 0.360 0.415 0.452
ETTh2 $\rightarrow$ ETTh1 96 0.585 0.510 0.541 0.503 0.732 0.577 1.130 0.777 0.485 0.465 0.689 0.555
192 0.677 0.554 0.559 0.515 0.758 0.559 1.242 0.820 0.565 0.509 0.707 0.568
336 0.700 0.562 0.620 0.551 0.759 0.578 1.328 0.864 0.581 0.515 0.710 0.577
720 0.693 0.579 0.729 0.627 0.781 0.597 4.145 1.461 0.628 0.561 0.704 0.596
Avg. 0.663 0.551 0.612 0.549 0.757 0.578 1.961 0.981 0.565 0.513 0.703 0.574
ETTh2 $\rightarrow$ ETTm2 96 0.181 0.288 0.218 0.304 0.253 0.329 0.646 0.563 0.226 0.309 0.240 0.336
192 0.235 0.324 0.265 0.335 0.293 0.346 0.934 0.654 0.289 0.345 0.295 0.369
336 0.294 0.357 0.327 0.370 0.347 0.376 1.157 0.728 0.348 0.379 0.345 0.397
720 0.395 0.411 0.401 0.416 0.446 0.429 4.730 1.531 0.439 0.427 0.432 0.442
Avg. 0.276 0.345 0.303 0.356 0.335 0.370 1.867 0.869 0.325 0.365 0.328 0.386
ETTm1 $\rightarrow$ ETTh2 96 0.415 0.437 0.331 0.383 0.353 0.392 0.510 0.576 0.354 0.385 0.365 0.415
192 0.486 0.477 0.353 0.399 0.443 0.437 0.523 0.586 0.447 0.434 0.454 0.462
336 0.397 0.433 0.400 0.428 0.469 0.461 0.640 0.637 0.481 0.463 0.496 0.494
720 0.451 0.475 0.417 0.448 0.466 0.468 2.296 1.034 0.474 0.471 0.541 0.529
Avg. 0.437 0.455 0.375 0.415 0.433 0.439 0.992 0.708 0.439 0.438 0.464 0.475
ETTm1 $\rightarrow$ ETTm2 96 0.081 0.185 0.194 0.270 0.217 0.294 0.646 0.563 0.195 0.271 0.221 0.314
192 0.162 0.250 0.243 0.304 0.277 0.327 0.934 0.654 0.258 0.311 0.286 0.359
336 0.250 0.311 0.295 0.341 0.331 0.360 1.157 0.728 0.317 0.348 0.357 0.406
720 0.376 0.392 0.367 0.385 0.429 0.413 4.730 1.531 0.416 0.404 0.476 0.476
Avg. 0.217 0.284 0.275 0.325 0.313 0.348 1.867 0.869 0.296 0.334 0.335 0.389
ETTm2 $\rightarrow$ ETTh2 96 0.485 0.473 0.322 0.369 0.360 0.401 0.510 0.576 0.327 0.367 0.333 0.391
192 0.496 0.481 0.359 0.396 0.434 0.437 0.523 0.586 0.411 0.418 0.441 0.456
336 0.542 0.509 0.439 0.452 0.460 0.459 0.640 0.637 0.439 0.447 0.505 0.503
720 0.437 0.463 0.448 0.468 0.485 0.477 2.296 1.034 0.459 0.470 0.543 0.534
Avg. 0.490 0.481 0.392 0.421 0.435 0.443 0.992 0.708 0.409 0.425 0.455 0.471
ETTm2 $\rightarrow$ ETTm1 96 0.345 0.377 0.446 0.415 0.747 0.558 1.179 0.781 0.491 0.437 0.570 0.490
192 0.518 0.462 0.496 0.452 0.781 0.560 1.327 0.846 0.530 0.470 0.590 0.506
336 0.735 0.541 0.507 0.463 0.778 0.578 1.478 0.902 0.565 0.497 0.706 0.567
720 0.653 0.533 0.556 0.482 0.769 0.573 3.749 1.408 0.686 0.565 0.731 0.584
Avg. 0.562 0.478 0.501 0.453 0.769 0.567 1.933 0.984 0.568 0.492 0.649 0.537